OpenStudy (anonymous):

?which point is the farthest from (2,-3

myininaya (myininaya):

\[(\lim_{n \rightarrow \infty} n,\lim_{p \rightarrow \infty} p)\]

myininaya (myininaya):

or negative infinity

myininaya (myininaya):

or +/- infinity

myininaya (myininaya):

i think your question is missing something though

OpenStudy (zarkon):

I would say there is no point that is farthest from (2,-3)

myininaya (myininaya):

lol

OpenStudy (anonymous):

Problem seems incomplete.

myininaya (myininaya):

zarkon why you gotta disagree but in a way it is like you aren't disagreeing because infinity doesn't exist lol

OpenStudy (anonymous):

Zarkon is being technically accurate. :)

myininaya (myininaya):

but i wanted to be crazy

OpenStudy (zarkon):

technically correct...the best kind of correct ;)

myininaya (myininaya):

but isn't giving crazy answers the best

myininaya (myininaya):

i think she was missing some likt which point is further from (2,-3) from the curve y=f(x)

OpenStudy (anonymous):

Or, a set of points.

myininaya (myininaya):

maybe that

OpenStudy (zarkon):

I like problems like that...closest point to a curve...fun stuff :)

myininaya (myininaya):

yes me too